OpenStudy (anonymous):

given that E=mc^2, find c if m=0.05 and E=4.5*10^15 (A) 1.5*10^7 (B) 3*10^8 (C) 1.0*10^15 (D) 2*10^16 or none of the above

OpenStudy (anonymous):

b

OpenStudy (anonymous):

i thought it was none of these can ou tell me how you got b so i know how to work it out?

OpenStudy (anonymous):

divide by m on both sides, u get E/m = c^2 E/m is 9 x 10^16 so c is sqrt{9 x 10^16} c=3 x 10^8

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

and anyway this c in this particular eqn is a universal constant

OpenStudy (anonymous):

what is a universal constant

OpenStudy (anonymous):

i mean this is one of the most famous equations of the world, and wherever u c dis eqn, d value of c is always the same

OpenStudy (anonymous):

thanks so much you have been a great help

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Errr, what kind of math is this? 0.0 -just stumbled upon this question-

OpenStudy (anonymous):

wts ur implication?